# SplatOverflow: Asynchronous Hardware Troubleshooting

### v2

@@ 10.1.2. Hardware Support for Maintainers
Through SplatOverflow, remote maintainers gain insight into the workspace and physical instance of the hardware a local user is working on. This was already possible with video conferencing and AR/VR workflows, but with SplatOverflow, maintainers can navigate a local user’s space independently. In SplatOverflow, maintainers do not rely on the local user to move a camera to inspect hardware. Maintainers can examine novel viewpoints independently and use gestures to guide the local user to move to the exact location.
This is a longstanding problem in collaborative workflows previously addressed with teleoperated robotic systems (Karchemsky et al., 2019; Sakashita et al., 2023). SplatOverflow achieves a similar end without the need for bespoke hardware or additional instrumentation of a space. This, paired with SplatOverflow’s web-based interface, expands the number of users who can access remote support.
@@ 10.1.3. Reducing Barriers to Distribute Hardware
Tech support is a core component of a viable hardware product but requires significant effort for hardware companies to scale. Asynchronous support infrastructure could reduce this effort by (a) providing communities of users the means of querying past issues and examining solutions and (b) allowing maintainers to provide support in an open format without coordinating with end-users. It would thus support maintainers, producers, and hardware communities in scaling technical support for users. To better evaluate this, we plan to engage with Open-source hardware producers and community maker spaces to study longer-term deployments of SplatOverflow.
@@ 10.2.1. Guided Splat Capture
In our evaluation, some user captures failed due to insufficient data to align the model. Given that proper alignment and registration are essential to a successful SplatOverflow scene, we plan to improve the feedback in SplatOverflow’s mobile interface. The capture process could be improved by specifying the hardware the user intends to scan. With this information, the interface can guide the user through capturing enough frames to appropriately align the splat to the CAD model and indicate which grounding and constraint tags require more footage.
@@ 10.2.3. IP of CAD models
In some cases, CAD models are highly protected by hardware maintainers, who may not want to share these files with users. Currently, SplatOverflow uses a mesh representation of the CAD model, and only needs to know where the alignment tags are placed. For maintainers who are concerned about sharing mesh models, we plan to build an import pipeline that strips only the relevant data for our workflows without saving the actual CAD model on our server.
@@ 10.2.4. Mixed Reality
10.3. Extending Multi-media workflows
Many tasks start with the assumption that the user is seeking guidance and has a screen available in front of them. Tools such as Interactive HTML BOM (Hackaday, 2018) allow users to visually correlate and search for components and their placements. The tool assumes users will be using a screen to aid in assembly and debugging.
Similarly, YouTube and other video hosting platforms (Utz and Wolfers, 2022; Kim et al., 2014; Bhatia, 2018) are used extensively for step-by-step instructions in cooking, assembly, and DIY Home repair. Documentation often has directly embedded video (Gibb, 2015, p. 121) and manufacturers such as IKEA (Wagstaff, 2012) have long produced videos to better demonstrate tasks. These videos can be valuable resources but are often expensive and challenging to create. As a result, video is not the ideal choice for end-users to articulate issues or seek guidance.
Our work extends multi-media hardware support by allowing users to rapidly author and share workflows for acting upon hardware amongst themselves. Communication can now more easily be made bidirectional rather than broadcast from experts or designers.
10.4. Cross-Referencing, Inspectability and Information Sharing
@@ 11. Conclusion
We have introduced SplatOverflow, a system for asynchronous hardware troubleshooting. SplatOverflow constructs a boundary object that can capture the physical details of a user’s hardware and communicate instructions as actions on the hardware instance. We demonstrate SplatOverflow through a series of workflows with different kinds of hardware. We have shown how these workflows can support complex troubleshooting workflows that require multiple exchanges between users and physical manipulation of the hardware. Moreover, we show how a SplatOverflow scene can be used in other application areas, such as querying data indexed to the CAD model and sharing maintenance and assembly workflows. Local users searching multiple disparate sources for relevant information can now access documentation and past solutions directly through the physical hardware. Maintainers relying on local users or telepresence robots to inspect remote spaces can now do so on their own, asynchronously.
We plan to distribute and deploy SplatOverflow with a broader user base through hardware maintainers and study our proposed forms of collaboration in the wild. Besides direct utility for hardware users and maintainers, we believe tools for communicating about physical hardware can reduce the barrier to entry for smaller producers distributing niche forms of hardware and supporting community development amongst their users.
@@ Acknowledgements.
We would like to thank the Digital Life Initiative at Cornell Tech for supporting this work through a doctoral fellowship. We would also like to thank Joey Castillo, Frank Bu and Stephen Hawes for taking part in preliminary discussions that helped motivate this work.References
@@
- Chen and Williams (1993) Shenchang Eric Chen and Lance Williams. 1993. View interpolation for image synthesis. In Proceedings of the 20th annual conference on Computer graphics and interactive techniques. ACM, Anaheim CA, 279–288. https://doi.org/10.1145/166117.166153
@@
- Houston et al. (2016) Lara Houston, Steven J. Jackson, Daniela K. Rosner, Syed Ishtiaque Ahmed, Meg Young, and Laewoo Kang. 2016. Values in Repair. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems. ACM, San Jose California USA, 1403–1414. https://doi.org/10.1145/2858036.2858470
@@
- Jackson and Kang (2014) Steven J. Jackson and Laewoo Kang. 2014. Breakdown, obsolescence and reuse: HCI and the art of repair. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. ACM, Toronto Ontario Canada, 449–458. https://doi.org/10.1145/2556288.2557332
@@
- McMillan and Bishop (1995) Leonard McMillan and Gary Bishop. 1995. An Image-Based Rendering System. Los Angeles (1995).
@@
- ORR (1996a) JULIAN E. ORR. 1996a. Talking about Machines: An Ethnography of a Modern Job. Cornell University Press. http://www.jstor.org/stable/10.7591/j.ctt1hhfnkz
- ORR (1996b) JULIAN E. ORR. 1996b. Talking about Machines: An Ethnography of a Modern Job. Cornell University Press. http://www.jstor.org/stable/10.7591/j.ctt1hhfnkz
- Penner and Zhang (2017) Eric Penner and Li Zhang. 2017. Soft 3D Reconstruction for View Synthesis. ACM Transactions on Graphics 36, 6 (Dec. 2017), 1–11. https://doi.org/10.1145/3130800.3130855
@@
- Wiles et al. (2020) Olivia Wiles, Georgia Gkioxari, Richard Szeliski, and Justin Johnson. 2020. SynSin: End-to-End View Synthesis From a Single Image. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, Seattle, WA, USA, 7465–7475. https://doi.org/10.1109/CVPR42600.2020.00749

### v3

@@ 9.1.2. Hardware Support for Maintainers
Through SplatOverflow, remote maintainers gain insight into the local user’s workspace and the actual state of the hardware they are working on. Moreover, maintainers do not have to rely on the local user to move a camera precisely to inspect hardware. They can examine novel viewpoints independently and use gestures to guide the local user to move to the exact location.
This is a longstanding problem in collaborative workflows previously addressed with teleoperated robotic systems (Karchemsky et al., 2019; Sakashita et al., 2023). SplatOverflow achieves a similar outcome without the need for bespoke hardware or additional instrumentation of a space. This, paired with SplatOverflow’s web-based interface, expands the number of users who can access remote support.
@@ 9.1.3. Reducing Barriers to Distribute Hardware
Tech support is a core component of a viable hardware product but requires significant effort for hardware companies to scale. Asynchronous support infrastructure could reduce this effort by (a) providing communities of users the means of querying past issues and examining solutions and (b) allowing maintainers to provide support in an open format without coordinating with end-users. This approach would support maintainers, producers, and hardware communities in scaling technical support for users. To better evaluate this, we plan to engage with open-source hardware producers and community maker spaces to study longer-term deployments of SplatOverflow.
9.2. Extending Multi-media Workflows
Many tasks start with the assumption that the user is seeking guidance and has a screen available in front of them. For example, tools such as Interactive HTML BOM (Hackaday, 2018) allow users to correlate components and their placements visually. The tool assumes users will be using a screen to aid in assembly and debugging.
Similarly, YouTube and other video hosting platforms (Utz and Wolfers, 2022; Kim et al., 2014; Bhatia, 2018) are used extensively for step-by-step instructions in cooking, assembly, and DIY Home repair. Documentation often has directly embedded video (Gibb, 2015, p. 121) and manufacturers such as IKEA (Wagstaff, 2012) produced videos to demonstrate operations clearly. While these videos can be valuable resources, they are often expensive and time-consuming to create. As a result, video is not always the ideal choice for end-users to articulate issues or seek guidance.
Our work extends multi-media hardware support by allowing users to rapidly author and share workflows for acting upon hardware amongst themselves. With SplatOverflow, bidirectional communication becomes easier instead of being solely broadcast from experts or designers.
9.3. Cross-Referencing, Inspectability and Information Sharing
@@ 9.4.1. Guided Splat Capture
In our evaluation, some user captures failed due to insufficient data to align the model. Given that proper alignment and registration are essential for creating a successful SplatOverflow scene, we plan to improve feedback in SplatOverflow’s mobile interface. The capture process could also be enhanced by allowing users to specify the hardware they intend to scan. With this information, the interface can guide users through capturing enough frames to align the splat to the CAD model and indicate which grounding and constraint tags require more footage.
@@ 9.4.3. IP of CAD models
In some cases, CAD models are highly protected by hardware maintainers, who may not want to share these files with users. Currently, SplatOverflow uses a mesh representation of the CAD model and only requires information about the position of alignment tags on the model. For maintainers concerned about sharing mesh models, we plan to build an import pipeline that strips only the relevant data for our workflows without saving the actual CAD model on our server.
@@ 9.4.4. Mixed Reality
10. Limitations
10.1. Acquiring CAD Models
SplatOverflow requires access to the hardware’s CAD model to populate the scene. For SplatOverflow scenes to be useful, the CAD models must also be complete. CAD models like this can be challenging to obtain outside of open-source hardware projects, which can limit the adoption of SplatOverflow. Moreover, even when CAD files are available, the structure of sub-assemblies, parts, and components is not standardized. As a result, preparing CAD files for use in SplatOverflow can require coercing the structure of the CAD assembly into a compatible shape.
10.2. Structure References in Technical Documentation
Another limitation of SplatOverflow lies in its ability to reference a corpus of existing technical documentation. With the Lumen v3, each technical documentation section contained links referencing the related parts. This greatly simplified populating the documentation interface with instructions to search through. However, not all hardware will share technical documentation in the same structured fields as Opulo does with Lumen v3. As a result, maintainers must post-process existing technical documentation to add the relevant structured data indicating which CAD components are referenced.
10.3. Errors in Tag Placement
@@ 11. Conclusion
We introduced SplatOverflow, a novel workflow for asynchronous hardware troubleshooting. SplatOverflow constructs a boundary object that can capture the physical details of a user’s hardware and communicate instructions as actions on their hardware. We demonstrate SplatOverflow through a series of examples with different kinds of hardware. We show how these workflows can support complex troubleshooting workflows that require multiple exchanges between users and physical manipulation of the hardware. Moreover, we illustrate how a SplatOverflow scene can index technical knowledge about hardware to the CAD model. As a result, local users searching multiple disparate sources for relevant information can now access documentation and past solutions directly through the physical hardware.
We plan to distribute and deploy SplatOverflow with a broader user base through hardware maintainers and study our proposed forms of collaboration in the wild. Besides direct utility for hardware users and maintainers, we believe tools for communicating about physical hardware can reduce barriers to entry for smaller producers distributing niche hardware and supporting community development amongst their users.
@@ Acknowledgements.
We thank the Digital Life Initiative at Cornell Tech for supporting this work through a doctoral fellowship. We thank the Bowers CIS Undergraduate Research Experience for supporting this work through their summer research program. We thank Roy Zunder for helping review this manuscript. Finally, we thank Joey Castillo, Frank Bu, and Stephen Hawes for participating in preliminary discussions that helped motivate this work.References
@@
- Bobrow and Whalen (2002) Daniel G. Bobrow and Jack Whalen. 2002. Community Knowledge Sharing in Practice: The Eureka Story. Reflections: The SoL Journal 4, 2 (Dec. 2002), 47–59. https://doi.org/10.1162/152417302762251336
@@
- Brand (2025) Stewart Brand. 2025. Maintenance: Of everything: Part one. Stripe Press.
@@
- Dunn et al. (2023) Kellie Dunn, Cynthia Feng, and Nadya Peek. 2023. Jubilee: A Case Study of Distributed Manufacturing in an Open Source Hardware Project. Journal of Open Hardware 7, 1 (May 2023), 4. https://doi.org/10.5334/joh.51
- Eghbal (2020) Nadia Eghbal. 2020. Working in public: the making and maintenance of open source software. Stripe Press.
@@
- Guédon and Lepetit (2023) Antoine Guédon and Vincent Lepetit. 2023. SuGaR: Surface-Aligned Gaussian Splatting for Efficient 3D Mesh Reconstruction and High-Quality Mesh Rendering. (2023). https://doi.org/10.48550/ARXIV.2311.12775 Publisher: [object Object] Version Number: 3.
@@
- Liu et al. (2024) Tianqi Liu, Guangcong Wang, Shoukang Hu, Liao Shen, Xinyi Ye, Yuhang Zang, Zhiguo Cao, Wei Li, and Ziwei Liu. 2024. MVSGaussian: Fast Generalizable Gaussian Splatting Reconstruction from Multi-View Stereo. http://arxiv.org/abs/2405.12218 arXiv:2405.12218.
@@
- ORR (1996) JULIAN E. ORR. 1996. Talking about Machines: An Ethnography of a Modern Job. Cornell University Press. http://www.jstor.org/stable/10.7591/j.ctt1hhfnkz
@@
- Tran O’Leary et al. (2024) Jasper Tran O’Leary, Thrisha Ramesh, Octi Zhang, and Nadya Peek. 2024. Tandem: Reproducible Digital Fabrication Workflows as Multimodal Programs. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–16. https://doi.org/10.1145/3613904.3642751
@@
- Wang et al. ([n. d.]) Shuzhe Wang, Vincent Leroy, Yohann Cabon, Boris Chidlovskii, and Jerome Revaud. [n. d.]. DUSt3R: Geometric 3D Vision Made Easy. ([n. d.]).